OpenStudy (anonymous):

need help!! How many solutions will the following system of linear equations have? –2x + 4y = –7 y=-1/2x+5 A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

1. you have to compare the slopes and the y intercepts. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. -2x + 4y = -7 4y = 2x - 7 y = 1/2x - 7/4 slope = 1/2 and y int = -7/4 y = -1/2x + 5 slope = -1/2 and y int = 5 if slopes and y intercepts are different,which in this problem they are, there is 1 solution.

OpenStudy (anonymous):

thanks for the help